OpenStudy (anonymous):

Find all singular points of the DE and determine whether each is regular or irregular x^2(x-5)^2y''+4xy'+(x^2-25)y=0

OpenStudy (anonymous):

\[y''+\frac{ 4 }{ x(x-5)^2 }y'+\frac{ x+5 }{ x^2(x-5) }y=0\]